### v2

@@ 3.3 Optimization Algorithm
The key to solving Problem (17) lies in leveraging the tensor structure to handle the HPSD constraint with reduced computational complexity.
Multiple Dual Subproblems: The proposed model in the original domain (Problem (17)) has a dual problem in the transform domain, since their solutions are uniquely connected via $\mathbf{M}$. As demonstrated below, the model is significantly easier to solve in the transform domain. First, we rewrite Problem (17) in the transform domain:
@@
$s.t.$ $\hat{\mathcal{A}}_{::i}\in H^{q}_{+},i\in[p],$
which can be decomposed into $p$ independent subproblems. We can divide and conquer each subproblem and gather the results to construct $\mathcal{A}$.
Derivative: For the $i$-th subproblem, we calculate the derivative concerning $\hat{\mathcal{A}}_{::i}$ and set the value to 0, yielding:
$\hat{\mathcal{A}}_{::i}=\left(\mathbf{S}_{i}-\frac{\eta_{i}}{2}\mathbf{I}_{q}\right)\left(\mathbf{S}_{i}+\lambda_{i}{\mathbf{W}_{i}}+\epsilon_{2}\mathbf{I}_{q}\right)^{-1},$ (19)
where $\mathbf{S_{i}}=\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}\left(\hat{\mathcal{X}}_{D^{o}}\right)^{T}_{::i}$ and $\mathbf{W}_{i}\in{\mathbb{R}^{q\times q}}$ is a diagonal matrix whose $j$-th diagonal element is calculated by:
$\mathbf{w}_{jj}=\left(1/\left(2\sqrt{\hat{\mathbf{a}}_{j}^{H}\hat{\mathbf{a}}_{j}+\epsilon_{1}}\right)\right).$ (20)
HPSD Projection: The next step is to project $\hat{\mathcal{A}}_{::i}$ onto the HPSD cone. This is achieved using the HPSD projection operator $P_{H^{d}_{+}}(\cdot)$ [55] which, for an arbitrary square matrix $\mathbf{A}\in\mathbb{C}^{b\times b}$, is defined as:
@@
$=\sum\limits_{i=1}^{b}max\left\{\sigma_{i},0\right\}\mathbf{u}_{i}^{H}\mathbf{u}_{i},$
where $\sum\nolimits_{i=1}^{b}\sigma_{i}\mathbf{u}_{i}^{H}\mathbf{u}_{i}$ is the eigenvalue decomposition of $\frac{1}{2}(\mathbf{A}+\mathbf{A}^{H})$.
Based on the above techniques, we solve the $i$-th subproblem of (18) with the following two-stage algorithm:
Step 1: Fix $\mathbf{W}_{i}$ and Update $\hat{\mathcal{A}}_{::i}$.
When $\mathbf{W}_{i}$ is fixed, the optimal $\hat{\mathcal{A}}_{::i}$ is obtained at:
@@
Step 2: Fix $\hat{\mathcal{A}}_{::i}$ and Update $\mathbf{W}_{i}$.
When $\hat{\mathcal{A}}_{::i}$ is fixed, $\mathbf{W}_{i}$ can computed directly by (20).
Discussion: Formulating a series of convex subproblems simplifies the optimization algorithm design. Although the HPSD projection requires eigenvalue decomposition with a computational complexity of $O(x^{3})$, the exploitation of tensor structures ensures that $x\ll\#{\rm Feature}$. The efficiency of the proposed method is demonstrated in the experimental results.
3.4 Feature selection
We assess the importance of an original feature by its contribution to the reconstruction. Given the presence of the transform domain, the influence of the transform matrix $\mathbf{M}$ must be considered.
@@ 3.4 Feature selection
which indicates that the contribution of $f_{lh}$ to the reconstruction of $f_{ij}$ can be described by $\sum^{p}_{k=1}m_{jh}m_{ik}a_{ilj}$. Therefore, the contribution matrix of $f_{lh}$ is defined as:
$\left[Contribution\right]_{ij}=\sum\limits^{p}_{l=1}\sum\limits^{q}_{h=1}\left(\sum\limits^{p}_{k=1}m_{jh}m_{ik}a_{ilj}\right).$ (24)
Finally, the score for $f_{lh}$ is computed as $\|Contribution\|^{2}_{F}$, which comprehensively captures its influence on the reconstruction of all features and leverages the global information. We summarize the whole algorithm in Algorithm 1.
Discussion: Equation (24) shows that an appropriate choice of $\mathbf{M}$ integrates feature correlation information from other modes into the mode targeted for slice-by-slice optimization. In our proposed method, each feature is scored by globally modeling its dependencies with other features without requiring sample vectorization, thus preserving the data organization and reducing computational complexity.
@@
0:  A third-order centralized data tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$, the order set $D^{o}$ determining $q$, an invertible matrices $\mathbf{M}\in\mathbb{C}^{q\times q}$, regularization parameters $\lambda,\eta$.
@@
1:  Randomly initialize $\left(\tilde{\mathcal{A}}\right)_{::i}\in{H^{q}_{+}}$, transform $\mathcal{X}$ into $\tilde{\mathcal{X}}$.
@@
3:     revolve $\tilde{\mathcal{X}}$ with $D^{o}$ and obtain $\tilde{\mathcal{X}}_{D^{o}}$.
4:     input $\left(\tilde{\mathcal{X}}_{D^{o}}\right)_{::i}$ and alternatively update $\hat{\mathcal{A}}_{::i}$ and $\mathbf{W}_{i}$ according to (22) until converge.
@@
6:  Compute $\mathcal{A}=\hat{\mathcal{A}}\times_{3}\mathbf{M}^{-1}$ and $\|Contribution\|^{2}_{F}$ for each score according to (24).
@@
8:  Sort $Score$ in descending order, and select features of the $h$ greatest scores. For slice-wise tensor, the score for each dimension will be obtained by summation.
@@ 4.1 Convergence analysis
First, we prove the convergence of the proposed algorithm. Let us introduce a lemma [40]:
@@
4.2 Computational complexity
Suppose $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ is rotated into $\mathcal{X}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$ ($n$ is the number of samples, $(p,q)\in\{(d_{1},d_{2}),(d_{2},d_{1})\}$). For each subproblem, it takes $O\left(q^{2}\right)$ to convert data into the transform domain and initialize the data covariance matrix, and $O\left(q^{3}\right)$ to update the reconstruction matrix in each iteration (including matrix inversion and HPSD projection). Therefore we have $p$ subproblems each solved at the computational cost of $O\left(q^{3}t_{i}+q^{2}n\right)$ ($t_{i}$ is the number of iterations for the $i$-th subproblem). The final feature score computation takes $O(p^{2}q)$. Eventually, the total computational complexity is $O\left(\sum\limits^{p}_{i=1}\left(q^{2}n+q^{3}t_{i}\right)+p^{2}q\right)$.
The number of features ($q\times p$) is usually much greater than the number of dimensions in a single mode ($q$ or $p$). Due to the exploitation of tensor structures with separate operations on each slice, STPCA-MP can handle data tensors at a remarkably smaller computation cost than non-tensor-based methods. For example, for image data of size $32\times 32$, non-tensor-based methods requiring matrix inversion or eigenvalue decomposition (such as SOGFS [40] and SPCAFS [38]; see Table 1) have a computational complexity of $O(1024^{3})$, whereas STPCA-MP requires only $O(32\times(32^{3}))$, reducing the complexity by $1024$ times. In addition, the computational complexity is linear to the number of samples, which means the proposed method will still be efficient when facing data tensors with numerous samples.
Table 1 shows the computational complexity of competitive methods in the following experiments. We can see that tensor-based methods reduce the computational burden from $\prod_{i}d_{i}^{x}$ to $d_{i}^{x}$ in each iteration by utilizing the tensor structure. Among tensor-based methods, CPUFS and CPUFSnn are the most recently proposed, which have a computational complexity linear to $d_{i}$ at the price of squared to $n$. They are non-convex models and usually take hundreds of iterations to converge, making it a challenge when facing large-scale datasets.
@@ 4.2 Computational complexity
SOGFS [40] matrix $O\left(dn^{2}+t(d^{3}t_{1}+n^{2}c)\right)$
RNE [50] matrix $O\left(dn^{2}+t(t_{1}(n+s)d^{2}+t_{2}n))\right)$
SPCAFS [39] matrix $O\left(d^{2}n+d^{3}t\right)$
NLGMS [56] matrix $O\left(t\left(d^{3}+n^{3}+nd+n^{2}d+d\log{d}\right)\right)$
CPUFS [18] tensor $O\left(d_{1}nc^{2}+d_{2}nc^{2}+d_{1}d_{2}nc+n^{2}c)t\right)$
CPUFSnn [18] tensor $O\left((d_{1}nc^{2}+d_{2}nc^{2}+d_{1}d_{2}nc+(n_{1}c)^{2}+(n_{2}c)^{2}+n^{2}c)t\right)$
MSPCA [33] tensor $O\left(t_{2}\sum_{i=1}^{2}(t_{1}+n)d_{i}^{3}\right)$
STPCA-MP tensor $O\left(\sum^{p}_{i=1}\left(q^{2}n+q^{3}t_{i}\right)+p^{2}q\right)(q\in\{d_{1},d_{2}\})$
@@ 5 Experiments
In this section, we conduct experiments on real-world data to demonstrate our proposed method’s effectiveness. The codes related to our proposed methods and the experiments are available at https://github.com/zjj20212035/STPCA.git.
@@ 5.1.1 Experimental environment
All the experiments are conducted on a personal computer with a 2.20 GHz CPU and 64GB main memory under the environment of the Windows 11 operation system.
@@ 5.1.2 Comparative methods
We compare our proposed methods with some state-of-the-art UFS methods as follows:
@@
    CPUFS and CPUFSnn [18] (2023): CPUFS is a tensor-based UFS method combining CP decomposition and spectral clustering. It utilizes a sparse classifier to perform UFS. CPUFSnn is a variation with nonnegative constraints on the classifier matrix.
@@
    MSPCA [33] (2014): MSPCA is a sparse tensor PCA method that utilizes Tucker decomposition to make use of tensor structure. It was originally proposed as a feature extraction method. We use the Kronecker product of its factor matrices to perform UFS, as an example of Tucker-based UFS methods.
@@
    NLGMS [56] (2024): NLGMS is a non-tensor-based method that introduces a self-representation technique into the graph-based framework to preserve global and local data structure. The self-representation term reconstructs each sample with the linear combination of other samples.
@@
    SPCAFS [38] (2021): SPCAFS is a non-tensor-based method that introduces $\ell_{2,p}$-norm into the one-dimensional PCA model. It forms a simple yet effective model.
@@
    RNE [50] (2020): RNE is a non-tensor-based method that makes use of the locally linear embedding algorithm to obtain an optimal projection matrix. It keeps the samples close to their neighbors after being projected onto a local manifold.
@@
    SOGFS [40] (2019): SOGFS is a non-tensor-based method that optimizes an adaptive graph to preserve the local geometrical structure of data. It updates the similarity matrix during each iteration to obtain a more flexible graph.
We also perform clustering with all features as a baseline method. All the comparative methods are implemented using the codes provided by their authors.

### v3

@@ 3.3 Optimization Algorithm
Solving Problem (18) efficiently requires leveraging the tensor structure to handle HPSD constraints with reduced computational complexity.
Multiple Dual Subproblems: Due to the bijective mapping $\hat{\mathcal{A}}=\mathcal{A}\times_{3}\mathbf{M}$, Problem (18) admits a dual formulation in the transform domain. This dual problem proves computationally tractable as shown below. We first recast it as:
@@
s.t. $\hat{\mathcal{A}}_{::i}\in H^{q}_{+},\quad i\in[p],$
which decomposes into $p$ independent subproblems. This enables a divide-and-conquer strategy: solve each frontal slice optimization separately, then reconstruct $\mathcal{A}$ post-optimization.
Derivative Analysis: For the $i$-th subproblem, setting the gradient with respect to $\hat{\mathcal{A}}_{::i}$ to zero yields:
$\hat{\mathcal{A}}_{::i}=\left(\mathbf{S}_{i}-\frac{\eta_{i}}{2}\mathbf{I}_{q}\right)\left(\mathbf{S}_{i}+\lambda_{i}{\mathbf{W}_{i}}+\epsilon_{1}\mathbf{I}_{q}\right)^{-1},$ (20)
where $\mathbf{S}_{i}=\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}\left(\hat{\mathcal{X}}_{D^{o}}\right)^{T}_{::i}$, $\epsilon_{1}>0$ is a small constant added for inversion stability, and $\mathbf{W}_{i}\in{\mathbb{R}^{q\times q}}$ denotes a diagonal matrix with $j$-th diagonal element:
$\mathbf{w}_{jj}=\left(1/\left(2\sqrt{\hat{\mathbf{a}}_{j}^{H}\hat{\mathbf{a}}_{j}+\epsilon_{2}}\right)\right).$ (21)
$\epsilon_{2}>0$ is a small constant to prevent the denominator from approaching zero.
HPSD Projection: Subsequently, $\hat{\mathcal{A}}_{::i}$ is projected onto the HPSD cone using the operator $P_{H^{d}_{+}}(\cdot)$ [58]. For an arbitrary square matrix $\mathbf{A}\in\mathbb{C}^{b\times b}$, this projection is defined through eigenvalue decomposition:
@@
$=\sum\limits_{i=1}^{b}\max\left\{\sigma_{i},0\right\}\mathbf{u}_{i}\mathbf{u}_{i}^{H},$
where $\frac{1}{2}(\mathbf{A}+\mathbf{A}^{H})=\sum\nolimits_{i=1}^{b}\sigma_{i}\mathbf{u}_{i}\mathbf{u}_{i}^{H}$ is its eigenvalue decomposition.
Building on these techniques, we solve the $i$-th subproblem of (19) using a two-stage iterative algorithm:
Step 1: Update $\hat{\mathcal{A}}_{::i}$ with $\mathbf{W}_{i}$ fixed. For fixed $\mathbf{W}_{i}$, the optimal $\hat{\mathcal{A}}_{::i}$ is given by:
@@
Step 2: Update $\mathbf{W}_{i}$ with $\hat{\mathcal{A}}_{::i}$ fixed. For fixed $\hat{\mathcal{A}}_{::i}$, $\mathbf{W}_{i}$ is directly computed via (21).
Discussion: This convex subproblem formulation significantly simplifies optimization design. While HPSD projection requires eigenvalue decomposition with $O(x^{3})$ complexity, tensor structure exploitation ensures $x\ll|\mathcal{F}|$ where $\mathcal{F}$ denotes the feature set. Experimental results demonstrate the method’s computational efficiency.
3.4 Feature selection via score map
We quantify feature importance by its contribution to tensor reconstruction, accounting for the transform domain induced by $\mathbf{M}$.
@@ 3.4 Feature selection via score map
The contribution of feature $f_{lh}$ to the reconstruction of all features is quantified by the matrix $\mathbf{C}^{lh}\in\mathbb{C}^{p\times q}$:
$\mathbf{C}^{lh}_{ij}=\sum_{k=1}^{p}m_{jh}m_{ik}a_{ilj}.$ (25)
The feature importance score is then defined as:
$score(f_{lh})=\|\mathbf{C}^{lh}\|_{F}^{2}=\sum_{i=1}^{p}\sum_{j=1}^{q}\left(\sum_{k=1}^{p}m_{jh}m_{ik}a_{ilj}\right)^{2},$ (26)
which comprehensively captures $f_{lh}$’s global influence on feature reconstruction. The complete procedure is summarized in Algorithm 1.
Discussion: Equation (25) demonstrates that an appropriately chosen $\mathbf{M}$ effectively integrates cross-mode feature correlations into the slice-by-slice optimization process. Our scoring mechanism globally quantifies feature dependencies through tensor operations, eliminating the need for sample vectorization. This approach provides dual advantages:
  • Preservation of intrinsic data organization structures
  • Significant reduction in computational complexity
by maintaining the native tensor representation throughout the feature evaluation process.
3.5 Implementation details
The proposed optimization framework decomposes into multiple independent subproblems amenable to parallel computation. We establish a unified convergence criterion for all subproblems:
$|\text{obj}^{(h)}-\text{obj}^{(h-1)}|<10^{-5},\quad h=1,2,3,\cdots$ (27)
where $\text{obj}^{(h)}$ denotes the objective function value at iteration $h$.
For regularization parameters $\lambda$ and $\eta$, we optimized them via grid search per dataset in the following comparison experiments. In practice, $\lambda$ and $\eta$ can be set to $10$ for a rapid deployment. The stabilization constants are fixed at $\epsilon_{1}=10^{-3}$ and $\epsilon_{2}=10^{-2}$ throughout all experiments.
For data preprocessing, we apply centralization to $\left(\mathcal{X}_{D^{o}}\right)_{::i}$ for each subproblem individually to capture inter-dimension variance more effectively. To narrow down the grid search range, we normalize $\mathcal{X}$ by dividing each element by the maximum absolute value, scaling all values to the interval [-1, 1].
@@ 3.5 Implementation details
0:  A third-order centralized data tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$, the order set $D^{o}$ determining $q$, an invertible matrices $\mathbf{M}\in\mathbb{C}^{q\times q}$, regularization parameters $\lambda,\eta$, and $\epsilon_{1}=10^{-3}$, $\epsilon_{2}=10^{-2}$.
@@
1:  Randomly initialize $\left(\hat{\mathcal{A}}\right)_{::i}\in{H^{q}_{+}}$, transform $\mathcal{X}$ into $\hat{\mathcal{X}}$.
@@
3:     revolve $\hat{\mathcal{X}}$ with $D^{o}$ and obtain $\hat{\mathcal{X}}_{D^{o}}$.
4:     input centralized $\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}$ and alternatively update $\hat{\mathcal{A}}_{::i}$ and $\mathbf{W}_{i}$ according to (23) until $|\text{obj}^{(h)}-\text{obj}^{(h-1)}|<10^{-5}$.
@@
6:  Compute $\mathcal{A}=\hat{\mathcal{A}}\times_{3}\mathbf{M}^{-1}$ and $\|\mathbf{C}^{lh}\|_{F}^{2}$ for each score according to (25).
@@
8:  Sort $score$ in descending order, and select features of the $h$ greatest scores. For a slice-wise tensor, the score for each dimension will be obtained by summation.
@@ 4.1 Convergence analysis
First, we prove the convergence of the proposed algorithm. The proof leverages the following fundamental lemma [43]:
@@
4.2 Computational complexity analysis
Consider a data tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ reoriented as $\mathcal{X}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$, where $n$ denotes the number of samples and $(p,q)\in\{(d_{1},d_{2}),(d_{2},d_{1})\}$. The computational cost comprises three main components:
  • Transform domain conversion: $O(q^{2})$ operations for data transformation and covariance matrix initialization per frontal slice.
  • Iterative optimization: $O(q^{3})$ per iteration for reconstruction matrix updates (including matrix inversion and HPSD projection).
  • Feature scoring: $O(p^{2}q)$ for final importance evaluation.
For $p$ independent subproblems, each requiring $t_{i}$ iterations, the total complexity is:
$O\left(\sum_{i=1}^{p}(q^{2}n+q^{3}t_{i})+p^{2}q\right)$ (38)
Computational Efficiency Advantages: Table 4 summarizes the computational complexity of comparable methods in subsequent experiments. The proposed STPCA-MP framework achieves significant computational efficiency through three core mechanisms: First, dimensionality reduction via tensor decomposition leverages the inherent structure by exploiting the inequality $q\times p\gg\max(q,p)$ through slice-by-slice operations, systematically replacing high-dimensional matrix computations with mode-decomposed equivalents and reducing per-iteration complexity from $\prod_{i}d_{i}^{x}$ to $d_{i}^{x}$. For $32\times 32$ image data, this reduces complexity from $O(1024^{3})$ (as in vectorized methods SOGFS [43] and SPCAFS [41]) to $O(32\times(32^{3}))$, representing a 1024-fold reduction. Second, linear sample scalability ensures complexity grows as $O(n)$, maintaining large-scale dataset efficiency, in contrast to the quadratic sample complexity $O(n^{2})$ incurred by graph-updating methods like CPUFS [21]. Third, the convex formulation with HPSD projection enables stable, rapid convergence, reducing iteration counts and further conserving runtime. Crucially, STPCA-MP uniquely maintains all advantages simultaneously, preserving the original tensor structure while achieving linear scaling in both feature dimensions ($d_{i}$) and sample size ($n$)—a dual efficiency unattained by approaches like CPUFS and CPUFSnn that sacrifice sample efficiency for dimensionality reduction.
@@ 4.2 Computational complexity analysis
SOGFS [^1] Structured Optimal Graph for Unsupervised Feature Selection [43] matrix $O\left(dn^{2}+t(d^{3}t_{1}+n^{2}c)\right)$
RNE [^2] Robust Neighborhood Embedding for unsupervised feature selection [53] matrix $O\left(dn^{2}+t(t_{1}(n+s)d^{2}+t_{2}n))\right)$
SPCAFS [^3] Sparse PCA via $\ell_{2,p}$-Norm regularization for unsupervised Feature Selection [42] matrix $O\left(d^{2}n+d^{3}t\right)$
NLGMS [^4] Unsupervised group feature selection approach via non-convex regularized graph embedding and self-representation [59] matrix $O\left(t\left(d^{3}+n^{3}+nd+n^{2}d+d\log{d}\right)\right)$
CPUFS [^5] CP decomposition based Unsupervised Feature Selection [21] tensor $O\left(d_{1}nc^{2}+d_{2}nc^{2}+d_{1}d_{2}nc+n^{2}c)t\right)$
CPUFSnn [^6] CP decomposition based Unsupervised Feature Selection with non-negativity [21] tensor $O\left((d_{1}nc^{2}+d_{2}nc^{2}+d_{1}d_{2}nc+(n_{1}c)^{2}+(n_{2}c)^{2}+n^{2}c)t\right)$
MSPCA [^7] Multilinear Sparse Principal Component Analysis [36] tensor $O\left(t_{2}\sum_{i=1}^{2}(t_{1}+n)d_{i}^{3}\right)$
STPCA-MP (ours) tensor $O\left(\sum^{p}_{i=1}\left(q^{2}n+q^{3}t_{i}\right)+p^{2}q\right)(q\in\{d_{1},d_{2}\})$
@@ 5 Experiments
In this section, we conduct experiments on real-world datasets to empirically validate the effectiveness of our proposed method. Implementation codes for STPCA-MP and all experiments are publicly hosted at: https://github.com/zjj20212035/STPCA.git.
@@ 5.1.1 Experimental environment
All experiments are executed on a standard workstation equipped with a 2.20 GHz CPU and 64 GB RAM, utilizing the Windows 11 operating system environment.
@@ 5.1.2 Comparative methods
We evaluate our method against representative state-of-the-art UFS baselines:
@@
    CPUFS and CPUFSnn [21] (2023): CPUFS is a tensor-based unsupervised feature selection method that integrates CP decomposition with spectral clustering. The method employs a sparse classifier to perform feature selection. CPUFSnn extends this framework by imposing non-negative constraints on the classifier matrix, enhancing interpretability through element-wise positivity requirements.
@@
    MSPCA [36] (2014): MSPCA represents a sparse tensor PCA approach employing Tucker decomposition to leverage inherent tensor structures. Originally developed as a feature extraction technique, we adapt it for unsupervised feature selection (UFS) by utilizing the Kronecker product of its factor matrices, thereby demonstrating a Tucker-based UFS implementation.
@@
    NLGMS [59] (2024): NLGMS constitutes a non-tensor-based approach that incorporates a self-representation mechanism within a graph-based framework to simultaneously maintain both global and local data structures. The method’s self-representation component reconstructs individual samples through linear combinations of other data instances.
@@
    SPCAFS [41] (2021): SPCAFS is a non-tensor-based approach that incorporates the $\ell_{2,p}$-norm regularization into the standard one-dimensional PCA framework. This formulation yields an efficient and computationally tractable model.
@@
    RNE [53] (2020): RNE represents a non-tensor-based approach that employs the locally linear embedding (LLE) algorithm to derive an optimal projection matrix. The method maintains neighborhood relationships by preserving proximity between samples and their local neighbors when projected onto the manifold space.
@@
    SOGFS [43] (2019): SOGFS is a non-tensor-based approach that optimizes an adaptive graph structure to maintain the local geometric relationships within the data. The method iteratively refines the similarity matrix to achieve progressively more flexible graph representations.
For baseline comparison, we additionally conduct clustering using the complete set of original features. All competing methods are implemented through their respective authors’ original code releases.